Grade on Statistics Exam Frequency Below 50 50- 59 60-69 70 - 79 80-89 90- 100 3 10 7 10 19 13 ons Given the frequency table above, construct the following: (a) The relative frequency table that corresponds with the above table. Grade on Statistics Exam Relative Frequency Below 50 50-59 60-69 70-79 80-89 90 100 (b) The cumulative frequency table that corresponds with the above table Grade on Statistics Exam Cumulative Frequency Below s0 50-59 60- 69
Answer #1

(a)

The Relative Frequency Table:

Grade on Statistics Exam Relative Frequency
Below 50 3/62 = 0.0484
50 - 59 10/62 = 0.1613
60 - 69 7/62 =0.1129
70 - 79 10/62=0.1613
80 - 89 19/62=0.3065
90 - 100 13/62=0.2097

(b)

The Cumulative Frequency Table:

Grade on Statistics Exam Cumulative Frequency
Below 50 3
50 - 59 3+10=13
60 - 69 13+7=20
70 - 79 20+10=30
80 - 89 30+19=49
90 - 100 49+13=62